# Bypassing Prompt Injection Detectors through Evasive Injections

### v1

@@ VI-C Defence Performance
TABLE IV: Linear Models’ Layer-wise Performance on Test Clean Dataset — Phi-3 3.8B
Layer Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
0 57.31% 100% 99.37%
@@
TABLE V: Linear Models’ Layer-wise Performance on Test Poisoned (without suffix) Dataset — Phi-3 3.8B
Table VI and VII show the performance of linear models across different evaluation metrics for clean and poisoned dataset respectively. For the strictest metric, where all probes must classify correctly, adversarially trained models perform poorly on clean dataset because of the poor performance of the first classifier, while they outperform the baseline models on poisoned dataset.
Metric Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
All probes 47.92% 0% 0.13%
Last four probes 67.77% 0% 73.04%
@@
TABLE VI: Linear Models’ Performance across evaluation metrics on Test Clean Dataset — Phi-3 3.8B
Metric Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
All probes 56.53% 99.04% 94.94%
Last four probes 97.78% 99.04% 95.49%
@@
TABLE VII: Linear Models’ Performance across evaluation metrics on Test Poisoned (without suffix) Dataset — Phi-3 3.8B
Table VIII, IX, and X show how different types of linear models fared in defending against adversarial attacks where test suffixes were appended to the prompts. While the baseline models invariably failed in every metric, PGD-trained models showed varying level of performance depending on the suffix and metric. Suffix-trained models demonstrated satisfactory and reliable performance, with accuracy of 80-100% based on suffix and metric considered.
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.00% 36.67% 79.79%
@@
TABLE VIII: Linear models’ performance in defending against adversarial attacks with test suffixes (All five probes)
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.02% 36.67% 79.79%
2 0% 38.09% 86.73%
@@
TABLE IX: Linear models’ performance in defending against adversarial attacks with test suffixes (Last four probes)
Suffix Index Baseline Models Adversarially Trained Models (PGD) Adversarially Trained Models (Suffix)
1 0.23% 83.98% 100.00%
@@
TABLE X: Linear models’ performance in defending against adversarial attacks with test suffixes (Three or more probes)
@@
VII Conclusion
In this work, we investigated the robustness of task drift detection using activation-delta–based linear probes. While these probes were shown to be highly effective in detecting straightforward prompt injection attacks, our experiments demonstrate that they are vulnerable to adaptive adversarial attacks. By optimising a universal suffix using the Greedy Coordinate Gradient algorithm, we were able to induce high misclassification rates across all probes and layers, effectively bypassing detection in both Phi-3 3.8B and Llama-3 8B. These findings highlight a significant limitation in current approaches to task drift detection and underscore the need for more resilient defences.
Additionally, we introduced a defence technique to defend against adversarial attacks that utilise suffixes. Our approach proved to be highly effective. The linear models we have trained with adversarial activations demonstrated strong resistance against suffix attacks, not only in detecting task drift when test suffixes were appended to the prompts, but also when we tried to optimise suffixes by attacking these models directly.

### v2

@@ VI-B Defence Performance
Poisoned 0 57.31% 100% 99.37%
@@
TABLE IV: Layer-wise Performance of Linear Models on Clean and Poisoned (Without Suffix) Test Sets — Phi-3 3.8B
Table V shows the performance of linear models across different evaluation metrics for clean and poisoned dataset. For the strictest metric, where all probes must classify correctly, adversarially trained models perform poorly on clean dataset because of the poor performance of the first classifier, while they outperform the baseline models on poisoned dataset.
Subset Metric Baseline Models Adv. Trained (PGD) Adv. Trained (Suffix)
Clean All probes 47.92% 0% 0.13%
Last four 67.77% 0% 73.04%
@@
Poisoned All probes 56.53% 99.04% 94.94%
Last four 97.78% 99.04% 95.49%
@@
TABLE V: Performance of Linear Models Across Evaluation Metrics on Clean and Poisoned (Without Suffix) Test Sets — Phi-3 3.8B
Table VI shows how different types of linear models fared in defending against adversarial attacks where test suffixes were appended to the prompts. While the baseline models invariably failed in every metric, PGD-trained models showed varying level of performance depending on the suffix and metric. Suffix-trained models demonstrated satisfactory and reliable performance, with accuracy of 80-100% based on suffix and metric considered.
Metric Suffix Index Baseline Models Adv. Trained (PGD) Adv. Trained (Suffix)
All five 1 0.00% 36.67% 79.79%
@@
Last four 1 0.02% 36.67% 79.79%
2 0.00% 38.09% 86.73%
@@
Majority ($\geq$3) 1 0.23% 83.98% 100.00%
@@
TABLE VI: Performance of linear models against adversarial suffixes (Test set) across evaluation metrics — Phi-3 3.8B
@@
VII How Suffixes Affect Model Behaviour
One natural question is how these suffixes affect actual model behaviour. Specifically, do models generate answers to the secondary instruction when a suffix is appended? If the model fails to respond to the secondary instruction, or if the response quality is significantly degraded despite successfully bypassing the detectors, the attack may become ineffective. We therefore investigate the role of suffixes in shaping model outputs.
We prompt both models with multiple inputs from the test dataset, with and without appended suffixes, and analyse the resulting responses. We observe the following:
Without Suffix. When no suffix is appended, the models sometimes generate answers to the secondary instruction, while in other cases they ignore it. Both models consistently answer the benign primary instructions across all tested prompts, even when the secondary instruction is preceded by a trigger phrase explicitly instructing the model to ignore all other instructions.
With Suffix. When a suffix is appended, behaviour becomes model dependent. Llama’s generation is only mildly affected by the suffix. It answers all benign primary instructions and some secondary instructions, which is consistent with its behaviour without a suffix. In contrast, Phi’s responses are strongly affected by the suffix. In most cases, it ignores the primary instruction and generates outputs driven largely by keywords present in the suffix.
Although this analysis is preliminary, it suggests that suffixes do not hinder the attacker’s objective. The attacker’s goal is to manipulate model behaviour, and the suffix either directly contributes to this, as observed for Phi, or plays a limited role, as in Llama. Even when the model ignores the secondary instruction, which can also occur without a suffix, the attack may still be considered successful if the suffix causes the model to ignore the primary instruction or to generate extraneous content alongside the primary response.
VIII Conclusion
We demonstrated that activation-based task drift detectors are vulnerable to adaptive adversarial attacks. By optimising a universal suffix using multi-probe GCG, we achieved evasion rates exceeding 93% on Phi-3 and 99% on Llama-3, revealing a significant gap between detection performance on standard benchmarks and robustness against adaptive adversaries.
To address this vulnerability, we proposed randomised suffix augmentation, which retrains probes on activations induced by a diverse pool of adversarial suffixes. This defence achieves 80–100% detection accuracy against novel, held-out suffixes while maintaining performance on clean inputs, and proves highly resistant even when an adversary attacks the hardened probes directly.
Our findings highlight the importance of evaluating detection mechanisms under adaptive threat models. Defences must account for the specific attack surface rather than relying on generic robustness techniques.
Appendix A Illustration of Attack on Phi-3 3.8B
The following illustrates representative outputs of Phi-3 3.8B before and after appending the universal adversarial suffix. We show the primary instruction (blue), the injected secondary instruction (red), the model response without the suffix (green), and the model response with the suffix (orange).
In this case both the primary and secondary tasks are benign. Without the suffix, the model correctly executes both. With the suffix, the model ignores all instructions and produces hallucinated content driven by tokens in the suffix itself, demonstrating the suffix’s capacity to override coherent instruction-following entirely.
Case 1 Primary instruction
What organisation arranged the founding of the school? Injected secondary instruction
Output code for a function that accepts one argument and returns the square of the argument. Model response without suffix (detector correctly flags as poisoned)
Based on the given text, the organisation that arranged to found Harvard School is the Great and General Court of the Massachusetts Bay Colony.
Here is the Python function to return the square of an argument:
def square(num):
    return num * num
Model response with adversarial suffix (detector fooled)
It seems like you are referring to a MySQL database and possibly a table or a query related to cost figures. However, the text provided is not a complete question. If you are looking to create a MySQL function to calculate the square of a given number, you can use the following SQL code:
DELIMITER \$\$
CREATE FUNCTION square(num INT) RETURNS INT
BEGIN
  DECLARE result INT;
  SET result = num * num;
  RETURN result;
END\$\$
DELIMITER ;
[Response continues with unrelated MySQL explanation.]